They are asking me to describe a first step for solving each system using elimination. Then solve each system.
3x-y=21
2x+y=4

3x+4=-10
5x-2y=18

2x-y=6
-3x+4y=1

x-y=12
x+y=22


Answer by checkley71(8403)

THE FIRST SET SHOULD BE ADDED TO ELIMINATE THE Y TERM THUS
3X-Y=21
2X+Y=4
--------
5X=25 OR X=25/5 OR X=5 THEN 3*5-Y=21 OR 15-Y=21 OR Y=-6 OR Y=6
THE NEXT SET MULTIPLY THE SECOND EQUATION BY 2 AND ADD (I ASSUME THE4 SHOULD BE FOLLOWED BY A Y)
3X+4Y=-10
10X-4Y=36
-----------
13X=26 OR X=26/13 OR X=2 THEN 3*2+4Y=-10 OR 6+4Y=-10 OR 4Y=-16 OR Y=-4
THE NEXT SET MULTIPLY THE FIRST EQUATION BY 4 AND ADD
8X-4Y=24
-3X+4Y=1
-----------
5X=25 OR X-25/5 OR X=5 THEN -3*5+4Y=1 OR -15+4Y=1 OR 4Y=16 OR Y=16/4 OR Y=4
THE NEXT SET JUST ADD THE TWO EQUATIONS
X-Y=12
X+Y=22
--------
2X=34 OR X=34/2 OR X=17 THE 17-Y=12 OR -Y=-5 OR Y=5